# Supernematic

### v1

@@ Abstract
Quantum theory of geometrically frustrated systems is usually approached as a gauge theory where the local conservation law becomes the Gauss law. Here we show that it can do something fundamentally different: enforce a global conserved quantity via a non-perturbative tiling invariant, rigorously linking microscopic geometry to a new macroscopically phase-coherent state. In a frustrated bosonic model on the honeycomb lattice in the cluster-charging regime at fractional filling, this mechanism protects a conserved global quantum number, the sublattice polarization $\tilde{N}=N_{A}-N_{B}$. Quantum fluctuation drives the spontaneous symmetry breaking of this global $U(1)$ symmetry to result in a supernematic phase — an incompressible yet phase-coherent quantum state that breaks rotational symmetry without forming a superfluid or realizing topological order. This establishes a route to a novel quantum many-body state driven by combinatorial constraints and boundary invariants. The realization of the SN on programmable quantum platforms can test ergodicity of the resonance operator in constrained quantum dynamics, and explore harnessing quantum hardware to explore classically hard tiling problems.
@@ I Introduction
Quantum systems at weak and strong coupling are governed by opposing organizing principles: kinetic energy favors delocalized, wave-like motion, while interaction energy drives particles toward optimal spatial organization. Although the uncertainty principle complicates descriptions that bridge these regimes, each limit is often simple to characterize. In the weak-coupling limit, the system forms a delocalized momentum eigenstate. In the strong-coupling limit, interactions can stabilize a unique spatial arrangement through crystallization, as in the Wigner crystal[1] or the Mott insulator[2]. Yet when additional internal degrees of freedom or geometric constraints frustrate these interactions, the optimization problem admits a manifold of equally favorable configurations rather than a single solution. In such cases, even the classical limit becomes profoundly nontrivial, giving rise to emergent gauge structures that connect degenerate states, as exemplified by spin ice[3] and dimer models[4]. Understanding how quantum fluctuations navigate this degenerate landscape—whether they select an ordered phase or preserve the imprint of frustration—remains one of the most conceptually challenging problems in modern quantum many-body physics.
@@
The fate of geometric frustration under quantum fluctuations in known cases is typically analyzed through gauge theories representing local constraints as Gauss laws. While deriving such continuum theories from generic lattice Hamiltonians is not analytically tractable, perturbing away from exactly soluble models has been productive [5, 6]. When the gauge theory is deconfined, the resulting state could be topologically ordered [7]. However, one may ask whether there is an alternative, rigorous path for the intricate microscopic geometrical constraints to leave their fingerprint on a quantum ground state. In this paper, we provide one such example. We will prove that geometrical frustration on the lattice can give rise to qualitatively new long-wavelength behavior that escapes a conventional gauge-theoretic description, without resorting to numerical simulations or uncontrolled approximations.
@@
Here we consider the effect of quantum fluctuations for bosons in a honeycomb lattice under cluster-charging interaction at commensurate fractional fillings (1/3, 2/3, 1, 4/3, and 5/3) as shown in Fig. 1 a) schematically. The strong coupling limit of this model, first inspired by the extended Wannier orbitals in twisted bilayer graphene[8], is equivalent to the problem of densely tiling triangles on faces of a triangular lattice[9]. Finding a tiling solution to a finite region with general tiles is a computationally hard problem, although we could leverage exact Bethe ansatz solution to predict the exponential degeneracy in the tiling solutions in this case.
In this paper, we study whether the degenerate configuration space of the complex optimization problem can impact the quantum ground state properties in the presence of finite quantum fluctuations. We find the powerful notion of tiling invariance by Conway and Lagrias [10] indicates an emergent quantum number for quantum dynamics within the constrained space of configurations. We predict new macroscopically coherent quantum ground state characterized by spontaneous breaking of the emergent $U(1)$ symmetry associated with the quantum number and discuss the properties of the new state. We describe the state as “supernematic” (SN), emphasizing that quantum fluctuations within the degenerate manifold favor anisotropic resonating units, whose spontaneous orientational alignment breaks the $C_{3}$ point-group symmetry.
@@
The rest of the manuscript is organized as follows. In Sec.II, we introduce the cluster charging model and discuss the smallest resonating operator which has anisotropic shape (See Fig.1b). In Sec.III, we prove the exactness of the emergent quantum number $\tilde{N}\equiv N_{A}-N_{B}$, the sublattice polarization, for the manifold satisfying cluster-charging constraints at a filling $2/3$ of the honeycomb lattice. We do this by mapping the configurations in the constrained Hilbert space to a tiling problem with a fixed tile set and using the Conway-Lagrias boundary invariant in Ref.[10]. In Sec. IV, we derive the effective field theory that can address the quantum phase transition associated with the spontaneous symmetry breaking of the emergent $U(1)$ symmetry. This is done through mapping the microscopic problem to a three-colored double-dimer model and then using the height-field representation of the dimer model. The resulting field theory describes a problem with intertwined orders with fields transform non-trivially under the emergent $U(1)$ symmetry, the lattice translation and the $C_{3}$ point group symmetry. We predict a continuous transition between a columnar-ordered phase (CO) that breaks translational symmetry and $C_{3}$ point-group symmetry (See Fig.1c), and a new macroscopically phase-coherent phase, which breaks the emergent $U(1)$ symmetry and the $C_{3}$ point-group symmetry, but not the lattice translation (See Fig.1d). Hence, we refer to this new phase as supernematic (SN). In Sec.V we discuss the properties of the columnar ordered phase and its melting via thermal fluctuations. In Sec.VI we discuss the properties of the supernematic phase and its thermal melting. The thermal phase diagram is summarized in Fig. 1e). In Sec.VIII, we summarize our results and discuss the implications.
@@ II Extended Hubbard model and the cluster charging regime
where the $n_{i}$ is summed over the sites of an hexagonal plaquette labelled by $r$, and the interaction is summed over all honeycomb plaquettes. Expanding $H_{U}$ in terms of density-density interaction, $4U=V_{1}=2V_{2}=2V_{3}$. This interaction is longer ranged than the on-site interaction of the Hubbard model, but it is not infinite-ranged like the Coulomb interaction. It has been shown that $H_{U}$ introduces non-trivial orbital geometrical frustration, resulting in extensively degenerate configurations and local $U(1)\times U(1)$ gauge symmetry [9, 11, 12] in the classical strong coupling limit. This intermediate-range interaction can arise when the interaction strengths in the general Hamiltonian Eq.1 follow a specific hierarchy. For the commensurate cases with filling $1$ and $2/3$ particles per unit cell, $H_{U}$ dominate when $\max\{V_{1},\frac{V_{2}}{2},\frac{V_{3}}{2}\}\gg|V_{1}-\frac{V_{2}}{2}|,|V_{1}-\frac{V_{3}}{2}|,|V_{2}-V_{3}|,V_{k},W$, where $k\geq 4$ denotes the further range interaction, and $W$ denotes the bandwidth. For filling $1/3$, $H_{U}$ dominate when $V_{1,2,3}\gg V_{k},W$.
@@
In Ref.[13], we showed that the classical ground state in the strong coupling limit of $H_{U}$ is the so-called “trimer liquid” state, which is an ensemble of equal probability of all the possible particle configurations that minimize $H_{U}$. The trimer liquid has power-law correlation functions and pinch-point structure factor of particle density. The question we would like to address is the fate of trimer liquid under bosonic quantum fluctuations in the cluster-charging regime. We will primarily focus on filling of $2/3$ particle per hexagonal plaquette and discuss generic commensurate fillings in Sec. VII.
@@
There are several differences between the cluster-charging regime of interest and two better-known strongly interacting quantum bosonic models: the Bose Hubbard model and the extended Bose Hubbard model. With the extremely short-ranged interaction of the Hubbard model, the strong coupling limit is atomic, where each site can host zero or one bosons, subject to the total fillings. Quantum fluctuation introduced through hopping drives superfluidity, resulting in the well-known phase diagram [14] with “Mott-lobes” anchored by integer fillings with superfluid phase in-between, in terms of chemical potential versus hopping strength. With the further-range interaction of the extended Bose Hubbard model, quantum fluctuation usually gives rise to superfluid or supersolid phases where both break the $U(1)$ symmetry associated with the total particle number. However, such phase coherence is forbidden for the cluster-charging Hamiltonian ground state since the total particle number is related to the sum of the cluster-charging over all the plaquettes, which are local gauge symmetries. By Elitzur’s theorem [15] the usual $U(1)$ symmetry no longer can be broken spontaneously, guaranteeing a correlated insulating state for our model. Now, the question is: what else can happen to the quantum ground state of the model?
@@
The first step for considering the effect of quantum fluctuation in geometrically frustrated problems is to identify the minimal resonating operator that connects different configurations and the symmetry these operators may preserve. In the canonical example of the dimer model, the effect of quantum fluctuations is wildly different between the bipartite and non-bipartite lattices. The local resonating operators admit a $U(1)$ gauge structure for bipartite lattice while a $Z_{2}$ gauge structure for non-bipartite lattice [16, 4]. Due to the confinement of the compact $U(1)$ gauge theory in $2+1D$ [17], the quantum dimer model on a bipartite lattice has gapped ground states that usually break lattice translational symmetry. On the other hand, the quantum dimer model on a non-bipartite lattice can have a gapped topologically ordered state, which is the deconfined phase of the $Z_{2}$ gauge theory [18]  [^1] A recently established powerful perspective for understanding the radical difference observes that the local resonating operator preserves a $U(1)$ 1-form symmetry associated wth the winding number of the dimers in the bipartite lattice, while only a $Z_{2}$ 1-form symmetry in the non-bipartite lattice. (see e.g. Ref.[40, 41] and the reference therein)
@@
In the cluster-charging regime, the local resonating operators depend on the filling fraction. For the filling of $2/3$, the operator involving the smallest number of nearest neighbor hopping is the operator we denote as $\mathcal{O}_{2}$ in Fig.1c), because a smaller resonating operator involving only one plaquette annihilates all the states in the constraint space (see Appendix Sec.C for the derivation). There are significant consequences to the extended local structure of the $\mathcal{O}_{2}$ taking the elongated shape (a similar scenario for filling $1/3$, which is called “lemniscate” in Ref. [12], also discussed in Sec.VII). Firstly, the elongated shape means that the resonant fluctuation can occur locally in three different orientations, $\alpha=1,2,3$. Secondly, the resonating operator $\mathcal{O}_{2}$ locally shuffles the sub-lattice polarization between the two neighboring plaquettes supporting the $\mathcal{O}_{2}$ operator. Hence, in the cluster-charging regime of filling $2/3$, the effective Hamiltonian can be written as
@@
We are interested in the phases of the quantum Hamiltonian $\mathcal{H}$ under finite quantum fluctuation $t_{2}$ and interactions. The emergent $U(1)\times U(1)$ gauge structure we found earlier in the classical limit by mapping the classical model to coupled dimer models (see Ref. [12] Appendix F, also see Sec.IV) carries over to the quantum Hamiltonian $\mathcal{H}$. If the local $U(1)\times U(1)$ symmetry were to be the only conservation law in the problem, one would expect quantum fluctuations to drive confinement and a gapped spectrum as in the quantum dimer model on a bipartite lattice. However, as we will show in the next section, the $\mathcal{O}_{2}$ operator globally preserves an additional quantum number: the sublattice polarization, $\tilde{N}\equiv N_{A}-N_{B}$, which will be proved to remain conserved under all local terms in $\mathcal{H}$. This conservation law is independent of the 1-form symmetry of the $U(1)\times U(1)$ gauge theory, indicating a richer structure beyond conventional confinement physics in quantum dimer models. The conservation of $\tilde{N}$ suggests a possible spontaneous symmetry breaking phase with macroscopic phase coherence, analogous to a superfluid, coexisting with the confinement of the $U(1)\times U(1)$ gauge theory: the supernematic (SN). As we show later, the SN has two features: the spontaneous symmetry breaking of $\tilde{N}$ conservation leads to gapless Goldstone mode and phase coherence, and the confinement of the $U(1)\times U(1)$ gauge field leads to the $C_{3}$ three-fold rotational symmetry breaking.
@@
Figure 1: a) The cluster charging model, with repulsive interaction upto third nearest neighbor. The blue shaded area denotes the range of the interaction. b) The smallest local move that changes the particle configuration locally, by applying the $\mathcal{O}_{2}$ operator, whose $\pm$ signs denote particle creation/annihilation. c) The colomnar order with alternating colomns of occupation on A sublattices (green) and B sublattices (orange). The unitcell is indicated by the red dashed lines. d) Top: the schematic wavefunction of the resonating state, with superposition of various tilings of the local resonating patterns (grey double-plaquette). Bottom: the grey double-plaquette denotes a local resonance of the two configurations. e) Schematic phase diagram of the quantum model, where $\tilde{t}/\tilde{\mathbf{y}}$ is the tuning parameter (see Sec.IV.3), and $T_{\eta}$ denotes the transition temperature of the $\vec{\eta}$-field, which is tuned by a different parameter of $\tilde{t}/\tilde{\mathbf{y}}$. Here the $T_{\eta}$ is picked such that all the thermal phases discussed in Sec.V.2 and Sec. VI.2 can be shown. For larger $T_{\eta}$, the area of PC and SL will shrink and these phases can disappear. TL: “trimer liquid”, which is the extensively degenerate classical ground state satisfying cluster-charging constraintes. CO: “colomnar ordering” (panel b)). PC: plaquette Coulomb. N: Nematic phase. SL: “Super”-fluid. SN: Supernematic (panel d)). TD: thermally disordered. The quantum phase transition between CO and SN at zero temperature is of 3D XY universality class. The blue lines denote the thermal phase transition of the three-state Potts universality class.

### v2

@@ Abstract
Quantum theory of geometrically frustrated systems is usually approached as a gauge theory where the local conservation law becomes the Gauss law. Here we show that it can do something fundamentally different: enforce a global conserved quantity via a non-perturbative tiling invariant, rigorously linking microscopic geometry to a new macroscopically phase-coherent state. In a frustrated bosonic model on the honeycomb lattice in the cluster-charging regime at fractional filling, this mechanism protects a conserved global quantum number, the sublattice polarization $\tilde{N}=N_{A}-N_{B}$. Quantum fluctuation drives the spontaneous symmetry breaking of this global $U(1)$ symmetry to result in a supernematic phase — an incompressible yet phase-coherent quantum state that breaks rotational symmetry without forming a superfluid or realizing topological order. This establishes a route to a novel quantum many-body state driven by combinatorial constraints. The realization of the SN on programmable quantum platforms can test the ergodicity of the resonance operator in constrained quantum dynamics and explore the use of quantum hardware to tackle classically hard tiling problems.
@@ I Introduction
Quantum systems at weak and strong coupling are governed by opposing organizing principles: kinetic energy favors delocalized, wave-like motion, while interaction energy drives particles toward optimal spatial organization. Although the uncertainty principle complicates descriptions that bridge these regimes, each limit is often simple to characterize. In the weak-coupling limit, the system forms a delocalized momentum eigenstate. In the strong-coupling limit, interactions can stabilize a unique spatial arrangement through crystallization, as in the Wigner crystal[Wigner] or the Mott insulator[Mott_1949]. Yet when additional internal degrees of freedom or geometric constraints frustrate these interactions, the optimization problem admits a manifold of equally favorable configurations rather than a single solution. In such cases, even the classical limit becomes profoundly nontrivial, giving rise to emergent gauge structures that connect degenerate states, as exemplified by spin ice[castelnovo2008magnetic] and dimer models[moessner2001short]. Understanding how quantum fluctuations navigate this degenerate landscape—whether they select an ordered phase or preserve the imprint of frustration—remains one of the most conceptually challenging problems in modern quantum many-body physics.
@@
The fate of geometric frustration under quantum fluctuations is often analyzed through gauge theories. The derivation of such theories are not analytically tractable in general, but perturbing away from exactly soluble models has been productive [RK, KITAEV20062]. When the gauge theory is deconfined, the resulting state could be topologically ordered [Moessner2001resonating]. However, one may ask whether there is an alternative, rigorous path by which the intricate microscopic geometric constraints imprint their signature on a quantum ground state. In this paper, we provide one such example. We will prove that geometrical frustration on the lattice can give rise to qualitatively new long-wavelength behavior that escapes a conventional gauge-theoretic description, without resorting to numerical simulations or uncontrolled approximations.
@@
Here we consider the effect of quantum fluctuations for bosons in a honeycomb lattice under cluster-charging interaction at commensurate fractional fillings (1/3, 2/3, 1, 4/3, and 5/3) as shown in Fig. 1 a) schematically. The strong coupling limit of this model, first inspired by the extended Wannier orbitals in twisted bilayer graphene[zhang2022fractional], is equivalent to the problem of densely tiling triangles on faces of a triangular lattice[trimer-bethe]. Finding a tiling solution for a finite region with general tiles is a computationally hard problem. However, we can leverage the exact Bethe ansatz to predict exponential degeneracy in the tiling solutions in this case.
In this paper, we study whether the degenerate configuration space of the complex optimization problem can impact the quantum ground state properties even amidst finite quantum fluctuations. We find the powerful notion of tiling invariance by Conway and Lagrias [conway1990tiling] to dictate the quantum dynamics within the constrained space of configurationsan to emergently conserve a global charge. This opens the possibility of a macroscopically coherent quantum ground state characterized by spontaneous breaking of an emergent $U(1)$ symmetry: the phase we term a “supernematic” (SN). The SN exhibits a new form of intertwined order, being phase-coherent (“super-”) while breaking $C_{3}$ rotational symmetry (“-nematic”), where the elongated shape of the resonating operator enables the latter. It is a rather unusual phase. It is an incompressible insulating state like a Mott insulator, hence not a superfluid. Yet it is a macroscopically phase-coherent state with gapless excitations.
@@
The rest of the manuscript is organized as follows. In Sec.II, we introduce the cluster charging model and discuss the smallest resonating operator with an elongated, lamniscate shape (See Fig.1b). In Sec.III, we prove the exactness of the emergent quantum number $\tilde{N}\equiv N_{A}-N_{B}$, the sublattice polarization, for the manifold satisfying cluster-charging constraints at a filling $2/3$ of the honeycomb lattice. We do this by mapping the configurations in the constrained Hilbert space to a tiling problem with a fixed tile set and using the Conway-Lagrias boundary invariant in Ref.[conway1990tiling]. In Sec. IV, we derive the effective field theory that can address the quantum phase transition associated with the spontaneous symmetry breaking of the emergent $U(1)$ symmetry. This is done through mapping the microscopic problem to a three-colored double-dimer model and then using the height-field representation of the dimer model. The resulting field theory describes a problem with intertwined orders with fields transform non-trivially under the emergent $U(1)$ symmetry, the lattice translation and the $C_{3}$ point group symmetry. We predict a continuous transition between a columnar-ordered phase (CO) that breaks translational symmetry and $C_{3}$ point-group symmetry (See Fig.1c), and a new macroscopically phase-coherent phase, which breaks the emergent $U(1)$ symmetry and the $C_{3}$ point-group symmetry, but not the lattice translation (See Fig.1d). Hence, we refer to this new phase as supernematic (SN). In Sec.V we discuss the properties of the columnar ordered phase and its melting via thermal fluctuations. In Sec.VI we discuss the properties of the supernematic phase and its thermal melting. The thermal phase diagram is summarized in Fig. 1e). In Sec.VIII, we summarize our results and discuss the implications.
@@ II Extended Hubbard model and the cluster charging regime
where the $n_{i}$ is summed over the sites of an hexagonal plaquette labelled by $r$, and the interaction is summed over all honeycomb plaquettes. Expanding $H_{U}$ in terms of density-density interaction, $4U=V_{1}=2V_{2}=2V_{3}$. This interaction is longer ranged than the on-site interaction of the Hubbard model, but it is not infinite-ranged like the Coulomb interaction. It has been shown that $H_{U}$ introduces non-trivial orbital geometrical frustration, resulting in extensively degenerate configurations and local $U(1)\times U(1)$ gauge symmetry [trimer-bethe, Giudice2022Phys.Rev.Ba, mao2023fractionalization] in the classical strong coupling limit. This intermediate-range interaction can arise when the interaction strengths in the general Hamiltonian Eq.1 follow a specific hierarchy. For the commensurate cases with filling $1$ and $2/3$ particles per unit cell, $H_{U}$ dominate when $\max\{V_{1},\frac{V_{2}}{2},\frac{V_{3}}{2}\}\gg|V_{1}-\frac{V_{2}}{2}|,|V_{1}-\frac{V_{3}}{2}|,|V_{2}-V_{3}|,V_{k},W$, where $k\geq 4$ denotes the further range interaction, and $W$ denotes the bandwidth. For filling $1/3$, $H_{U}$ dominate when $V_{1,2,3}\gg V_{k},W$.
@@
In Ref.[zhang2024bionicfractionalizationtrimermodel], we showed that the classical ground state in the strong coupling limit of $H_{U}$ is the so-called “trimer liquid” state, which is an ensemble of equal probability of all the possible particle configurations that minimize $H_{U}$. The trimer liquid has power-law correlation functions and pinch-point structure factor of particle density. The question we would like to address is the fate of trimer liquid under bosonic quantum fluctuations in the cluster-charging regime. We will primarily focus on filling of $2/3$ particle per hexagonal plaquette and discuss generic commensurate fillings in Sec. VII.
@@
There are several differences between the cluster-charging regime of interest and two better-known strongly interacting quantum bosonic models: the Bose Hubbard model and the extended Bose Hubbard model. With the extremely short-ranged interaction of the Hubbard model, the strong coupling limit is atomic, where each site can host zero or one bosons, subject to the total fillings. Quantum fluctuation introduced through hopping drives superfluidity, resulting in the well-known phase diagram [Fisher1989Phys.Rev.B] with “Mott-lobes” anchored by integer fillings with superfluid phase in-between, in terms of chemical potential versus hopping strength. With the further-range interaction of the extended Bose Hubbard model, quantum fluctuation usually gives rise to superfluid or supersolid phases where both break the $U(1)$ symmetry associated with the total particle number. However, such phase coherence is forbidden for the cluster-charging Hamiltonian ground state since the total particle number is related to the sum of the cluster-charging over all the plaquettes, which are local gauge symmetries. By Elitzur’s theorem [elitzur1975impossibility] the usual $U(1)$ symmetry no longer can be broken spontaneously, guaranteeing a correlated insulating state for our model. Now, the question is: what else can happen to the quantum ground state of the model?
@@
The first step for considering the effect of quantum fluctuation in geometrically frustrated problems is to identify the minimal resonating operator that connects different configurations and the symmetry these operators may preserve. In the canonical example of the dimer model, the effect of quantum fluctuations is wildly different between the bipartite and non-bipartite lattices. The local resonating operators admit a $U(1)$ gauge structure for bipartite lattice while a $Z_{2}$ gauge structure for non-bipartite lattice [SubirKagome, moessner2001short]. Due to the confinement of the compact $U(1)$ gauge theory in $2+1D$ [POLYAKOV1977429], the quantum dimer model on a bipartite lattice has gapped ground states that usually break lattice translational symmetry. On the other hand, the quantum dimer model on a non-bipartite lattice can have a gapped topologically ordered state, which is the deconfined phase of the $Z_{2}$ gauge theory [FS79]  [^1] A recently established powerful perspective for understanding the radical difference observes that the local resonating operator preserves a $U(1)$ 1-form symmetry associated wth the winding number of the dimers in the bipartite lattice, while only a $Z_{2}$ 1-form symmetry in the non-bipartite lattice. (see e.g. Ref.[higherformLSM, Pace2023Phys.Rev.B] and the reference therein)
@@
In the cluster-charging regime, the local resonating operators depend on the filling fraction. For the filling of $2/3$, the operator involving the smallest number of nearest neighbor hopping is the operator we denote as $\mathcal{O}_{2}$ in Fig.1c), because a smaller resonating operator involving only one plaquette annihilates all the states in the constraint space (see Appendix Sec.C for the derivation). There are significant consequences to the extended local structure of the $\mathcal{O}_{2}$ taking the elongated shape (a similar scenario for filling $1/3$, which is called “lemniscate” in Ref. [mao2023fractionalization], also discussed in Sec.VII). Firstly, the elongated shape means that the resonant fluctuation can occur locally in three different orientations, $\alpha=1,2,3$. Secondly, the resonating operator $\mathcal{O}_{2}$ locally shuffles the sub-lattice polarization between the two neighboring plaquettes supporting the $\mathcal{O}_{2}$ operator. Hence, in the cluster-charging regime of filling $2/3$, the effective Hamiltonian can be written as
@@
We are interested in the phases of the quantum Hamiltonian $\mathcal{H}$ under finite quantum fluctuation $t_{2}$ and interactions. The emergent $U(1)\times U(1)$ gauge structure we found earlier in the classical limit by mapping the classical model to coupled dimer models (see Ref. [mao2023fractionalization] Appendix F, also see Sec.IV) carries over to the quantum Hamiltonian $\mathcal{H}$. If the local $U(1)\times U(1)$ symmetry were to be the only conservation law in the problem, one would expect quantum fluctuations to drive confinement and a gapped spectrum as in the quantum dimer model on a bipartite lattice. However, as we will show in the next section, the $\mathcal{O}_{2}$ operator globally preserves an additional quantum number: the sublattice polarization, $\tilde{N}\equiv N_{A}-N_{B}$, which will be proved to remain conserved under all local terms in $\mathcal{H}$. This conservation law is independent of the 1-form symmetry of the $U(1)\times U(1)$ gauge theory, indicating a richer structure beyond conventional confinement physics in quantum dimer models. The conservation of $\tilde{N}$ suggests a possible spontaneous symmetry breaking phase with macroscopic phase coherence, analogous to a superfluid, coexisting with the confinement of the $U(1)\times U(1)$ gauge theory: the supernematic (SN). As we show later, the SN has two features: the spontaneous symmetry breaking of $\tilde{N}$ conservation leads to gapless Goldstone mode and phase coherence, and the confinement of the $U(1)\times U(1)$ gauge field leads to the $C_{3}$ three-fold rotational symmetry breaking.
@@
Figure 1: a) The cluster charging model, with repulsive interaction upto third nearest neighbor. The blue shaded area denotes the range of the interaction. b) The smallest local move that changes the particle configuration locally, by applying the $\mathcal{O}_{2}$ operator, whose $\pm$ signs denote particle creation/annihilation. c) The colomnar order with alternating colomns of occupation on A sublattices (green) and B sublattices (orange). The unitcell is indicated by the red dashed lines. d) Top: the schematic wavefunction of the resonating state, with superposition of various tilings of the local resonating patterns (grey double-plaquette). Bottom: the grey double-plaquette denotes a local resonance of the two configurations. e) Schematic phase diagram of the quantum model, where $\tilde{t}/\tilde{\mathbf{y}}$ is the tuning parameter (see Sec.IV.3), and $T_{\eta}$ denotes the transition temperature of the $\vec{\eta}$-field, which is tuned by a different parameter of $\tilde{t}/\tilde{\mathbf{y}}$. Here the $T_{\eta}$ is picked such that all the thermal phases discussed in Sec.V.2 and Sec. VI.2 can be shown. For larger $T_{\eta}$, the area of PC and SL will shrink and these phases can disappear. TL: “trimer liquid”, which is the extensively degenerate classical ground state satisfying cluster-charging constraintes. CO: “colomnar ordering” (panel b)). PC: plaquette Coulomb. N: Nematic phase. SL: “Super”-fluid. SN: Supernematic (panel d)). TD: thermally disordered. The quantum phase transition between CO and SN at zero temperature is proposed to be of 3D XY universality class. The blue lines denote the thermal phase transition of the three-state Potts universality class.